# PSST: A Benchmark for Evaluation-driven Text Public-Speaking Style Transfer

### v2

@@ 1 Introduction
Benefiting from their superior capabilities, large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2023) can achieve high performance in traditional style transfer tasks. However, traditional style transfer is limited to sentence-level transformations and only a few data-driven styles (Jin et al., 2022) (e.g., formal-informal (Sheikha and Inkpen, 2011), polite-impolite (Madaan et al., 2020)). This is insufficient for complex real-world applications, which require passage-level transformations and complex styles (Jin et al., 2022). In this paper, we introduce a novel task, called Public-Speaking Style Transfer (PSST), which could further facilitate critical applications such as knowledge dissemination, public education, business promotion, etc. Specifically, PSST transforms formal, long paragraph texts into a public-speaking style, a form of language used by human beings to convey influential knowledge and ideas in public (Kedrowicz and Taylor, 2016; Beebe and Beebe, 2005) (as shown in Figure 1 and video demo in F.1).
@@
To investigate the language style modeling capabilities of LLMs in scenarios akin to PSST, we initially employ LLMs (eg. ChatGPT OpenAI (2022) and Llama 3-Instruct (AI@Meta, 2024)) for PSST and identify three primary issues: (1). Over-stylization; (2). Uneven Style Strength Distribution; (3). Severe Semantic Degradation. To further quantify and enhance the language style modeling capabilities of LLMs, we propose a fine-grained evaluation framework tailored for complex long-text style transfer tasks such as PSST focusing on style strength and semantic preservation (Rao and Tetreault, 2018; Shen et al., 2017a), which enables a continuous, evaluation-driven approach to enhance the language style modeling capabilities of LLMs.
@@
For style strength evaluation, we propose two metrics: (1) passage-level style strength score that offers a coarse-grained measure of the overall style strength, and (2) style strength distribution that captures the distribution of style elements throughout the text. Using real data as a standard, we analyze the style modeling capability of LLMs based on the above metrics. Specifically, we first gather official texts (eg. Encyclopedias) as source texts, and real data (eg. Ted Talks) as the target style dataset. To accurately define and evaluate public-speaking style, we review linguistic literature on spoken public speeches (McCroskey et al., 2003; Beebe and Beebe, 2005; Atkinson, 1985; Halliday, 1989) and summarize four prominent sub-styles, including orality, interactivity, vividness, and emotionality, supported by manual annotation. Since we aim to quantify the style elements at various positions within the stylized text and LLMs possess strong capabilities at the sentence level TST (Lai et al., 2023), for each sub-style, we generate a series of examples varying in style strength from a single sentence, and then we rank and score them with GPT3.5-Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo and train a small scorer (TinyLlama-1.1b (Zhang et al., 2024)) to predict these scores.
@@
For semantic preservation evaluation, we propose a QA-based method. The accuracy changes of a QA model (eg. Llama3-8B-Instruct (AI@Meta, 2024)) on texts before and after style transfer reflect the LLM’s ability of semantic preservation. Specifically, we use GPT-4333https://openai.com/research/gpt-4 to generate high-quality QA pairs from the source text focusing on two dimensions: key message and logical structure (Coopman and Lull, 2018).
@@
  • We collect data and introduce a valuable and extensible task named Public-Speaking Style Transfer (Section 3). We decompose the intricate text style into key sub-styles for more accurate style definition and evaluation (Section 3.3 3.4).
@@
Table 1: Comparison of PSST and formality TST: a. ’p-s’ means public-speaking style. b. Traditional TST primarily focuses on vocabulary-level adjustments, such as adhering to writing norms and word norms. In contrast, PSST involves audience engagement, including posing questions to generate interest and appeal to the audience.
@@ 2 Related Work
Style, from a linguistic perspective, encompasses various elements that contribute to the conveyance of semantics, including word choice, sentence structure, and arrangement, all of which work together to establish the tone, imagery, and meaning in the text (McDonald and Pustejovsky, 1985; Hu et al., 2022). In contrast, research on text style transfer (TST) takes a data-driven approach, defining style as attributes or labels based on style-specific corpora  (Shen et al., 2017b; Rao and Tetreault, 2018), which may affected by other attributes in dataset (Jin et al., 2022). However, our approach is grounded in real data and explores key dimensions from the perspective of public-speaking linguistics. This definition alleviates the ambiguity and reduces the difficulty of data construction and evaluation.
@@
Firstly, fluency, a common objective in most natural language generation tasks, is often measured by the perplexity score (PPL) (Yang et al., 2018). Secondly, to evaluate content preservation during the style transfer, metrics include BLEU (Papineni et al., 2002), ROUGE (Lin and Och, 2004), BERTScore (Zhang et al., 2019) are employed. Thirdly, the strength of style is an important dimension. Typically, a binary style classifier is first separately pretrained to predict the style label of input sentences (Prabhumoye et al., 2018; Lai et al., 2021; Zhan et al., 2022). This classifier is then used to estimate the style transfer accuracy. Recently,  Lai et al. (2023) demonstrates that ChatGPT (OpenAI, 2022) achieves competitive correlation with human judgments to serve as a multidimensional evaluator for sentence-level formal-informal style transfer.
@@ 3.4 Key Features of Public-Speaking Style
  • Interactivity: Interactivity in public speaking refers to the speaker engaging with the audience through various means such as posing thought-provoking, facilitating personal reflection, and crafting intriguing hypothetical scenarios. (Table 1 example 1)
  • Emotionality: Public speaking contains the speaker’s appropriate views and attitudes on specific events to reflect the speaker’s emotional tendencies and inner thoughts. (Table 1 example 2)
  • Vividness: In public speaking, speakers should present information in a lively, easy-to-understand way, such as using analogies and metaphors to make complex ideas more accessible and engaging. (Table 1 example 3)
  • Orality: The public-speaking style texts should align with oral communication norms, incorporating appropriate filler words, simple sentence structures, and suitable word choices.(Table 1 example 4)
@@ 4.1 Style Strength Evaluation
To evaluate the style strength of long texts in the PSST task, we propose a multi-dimensional, fine-grained evaluation method. Specifically, we generate a series of examples with varying style strengths from a single sentence, rank and score them using gpt-3.5-turbo, and then train a small scorer (TinyLlama-1.1b (Zhang et al., 2024)) to predict these scores. This approach is appropriate for three main reasons: 1). Early experiments suggest that using LLMs directly as evaluators may not be feasible (Section 4.1.3). 2). Relevant studies show that LLMs possess strong stylization and evaluation capabilities in TST tasks at the sentence level (Lai et al., 2023). 3). For long texts, a simple overall style strength score does not facilitate detailed analysis. We believe that a similar style of two long texts should be reflected in the overall style strength and the stylization of specific positions within the documents, such as the beginning and the end.
@@ 4.2.2 Correlation
We conduct three manual evaluation experiments to verify the effectiveness of the QA-based method. Please refer to Appendix E.3 for more details.
@@
The experiments above demonstrate that the decreased accuracy of the QA model is indeed attributed to the semantic loss occurring during the style transfer process.
@@ 5.1 Experiment Settings
We assessed the public-speaking style modeling capability of the LLM by comparing it with the following text. (1). ST(source text), the official texts mentioned in Section 3.2, characterized by an official style and complete information, serving as the upper bound for semantic preservation. (2). PT(Paraphrased text), texts paraphrased by gpt-3.5-turbo based on the source text, reflecting inherent changes in the language processing of LLMs. (3). Ted Talks, the target style text in real public speaking scenario. Note that ST and TED are not parallel. Both ST and PT have low public-speaking style strength but high semantic preservation scores. The closer the style strength of the generated texts is to real-world data and the lower their semantic loss, the stronger the model’s capability in language style modeling.
@@
(1). concise prompt, a brief instruction requiring models to simulate lively and engaging public speaking through style transfer. (2). enhanced prompt, detailed guidance on the dimensions of stylization, cautioning against excessive stylization, and emphasizing the importance of semantic consistency. Five prompts are created for each type shown in Figure 16.
@@
We employ Llama 2-Chat (Touvron et al., 2023), Llama 3-Instruct (AI@Meta, 2024) and gpt-3.5-turbo to generate public-speaking style texts based on ST and Prompts. We believe these models capture the current state of LLMs and can help explore the gap between LLMs and human abilities in complex, abstract language style modeling.
@@ 5.2 Evaluation Results and Analysis
The results comparing style strength distribution across different positions within texts are illustrated in Figure 5. Specifically, in terms of interactivity and orality, the style distribution in real-world scenarios is more uniform, with slight increases in stylization at the opening and closing segments. This pattern aligns with the conventional use of introductory and concluding phrases in real-world settings. However, for LLMs, the distribution exhibits a U-shaped pattern, indicating excessive stylization at the text’s beginning and end. For instance, as depicted in Figure 7, compared to real-world texts, LLMs such as GPT-3.5, Llama-3-70B-Instruct, and Llama-2-Chat-70B demonstrate pronounced stylization at the beginning and end, but less so in the middle sections(Case in Figure 21). This disparity suggests that LLMs diverge from human-like style modeling, potentially due to a model-inherent tendency to focus on the beginning and end of long texts, or possibly indicative of the model’s ’lazy’ behavior.
@@
The results shown in Figure 6 reveal significant semantic loss during style transfer. The high accuracy observed in src_text and paraphrase confirms that the QA model can effectively respond to questions based on non-stylized texts. Moreover, our questions, designed to target essential information, are relatively straightforward and can be answered provided no information is lost. Therefore, the observed decline in accuracy with stylized texts further underscores the challenges LLMs face in semantic preservation(Case 3, 4 in Figure 2425). Notably, models exhibiting stronger stylization (e.g., LLaMA-3-70B-Instruct vs. GPT-3.5-Turbo) tend to obscure key messages more. This is intuitive, as excessive stylization reduces the clarity of semantic content, thus hindering message comprehension for audiences, even if the essential points remain present. Additionally, while enhanced prompts focus on semantic preservation and achieve higher accuracy, they are still less than satisfactory.
@@ 6 Conclusion
In this paper, we introduce public-speaking style transfer, which requires LLMs to transform a formal and long text into a public-speaking style. By analyzing real-world data from a linguistic perspective, decompose the intricate text style into key sub-styles. We propose a fine-grained evaluation framework that enables an evaluation-driven approach to continuously assess and enhance the language style modeling capabilities of LLMs. For style strength evaluation, we distill the LLM’s ability to assess the style elements of sentences into a smaller model for each sub-style. We introduce two metrics: paragraph-level stylization intensity score and stylization intensity distribution, allowing for a detailed analysis of stylized texts. For semantic preservation, we propose a QA-based method that uses QA pairs focusing on key information and logical structure. By observing changes in the QA model’s accuracy, we identify and analyze the LLM’s ability to retain semantics. Our experiments reveal that current LLMs exhibit issues such as over-stylization, uneven style strength distribution, and severe semantic degradation in long and complicated language style modeling.
@@ Acknowledgements
We thank Xiyue Liang and Shunyu Wang for their support in human annotation for this study.
@@ C.1 Results of Different Text Lengths in PSST
Table 7: Style Strength Evaluation Results($\textbf{800}\pm 200$ tokens). a. The original score ranges from 0 to 5. b. ST means Source Text and PT means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).
@@
Table 8: Style Strength Evaluation Results($\textbf{400}\pm 100$ tokens). a. The original score ranges from 0 to 5. b. ST means Source Text and PT means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).
@@
Table 10: Style Strength Evaluation Results($\textbf{1200}\pm 100$ tokens). a. The original score ranges from 0 to 5. b. ST means Source Text and PT means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).

### v3

@@ 1 Introduction
Benefiting from their superior capabilities, large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2023) can achieve high performance in traditional style transfer tasks (Li et al., 2024a). However, traditional style transfer is limited to sentence-level transformations and only a few data-driven styles (Jin et al., 2022) (e.g., formal-informal (Sheikha and Inkpen, 2011), polite-impolite (Madaan et al., 2020)). This is insufficient for complex real-world applications, which require passage-level transformations and complex styles (Jin et al., 2022). In this paper, we introduce a novel task, called Public-Speaking Style Transfer (PSST), which could further facilitate critical applications such as knowledge dissemination, public education, business promotion, etc. Specifically, PSST transforms formal, long paragraph texts into a public-speaking style, a form of language used by human beings to convey influential knowledge and ideas in public (Kedrowicz and Taylor, 2016; Beebe and Beebe, 2005; Gao et al., 2020) (as shown in Figure 1 and video demo in F.1).
@@
To investigate the language style modeling capabilities of LLMs in scenarios akin to PSST, we initially employ LLMs (eg. ChatGPT OpenAI (2022) and Llama 3-Instruct (AI@Meta, 2024)) for PSST and identify three primary issues: (1) Over-stylization; (2) Uneven Style Strength Distribution; (3) Severe Semantic Degradation. To further quantify and enhance the language style modeling capabilities of LLMs, we propose a fine-grained evaluation framework tailored for complex long-text style transfer tasks such as PSST focusing on style strength and semantic preservation (Rao and Tetreault, 2018; Shen et al., 2017a), which enables a continuous, evaluation-driven approach to enhance the language style modeling capabilities of LLMs.
@@
For style strength evaluation, we propose two metrics: (1) passage-level style strength score that offers a coarse-grained measure of the overall style strength, and (2) style strength distribution that captures the distribution of style elements throughout the text. Using real data as a standard, we analyze the style modeling capability of LLMs based on the above metrics. Specifically, we first gather official texts (eg. Encyclopedias) as source texts, and real public speaking data (eg. Ted Talks) as the target style dataset. To accurately define and evaluate public-speaking style, we review linguistic literature on spoken public speeches (McCroskey et al., 2003; Beebe and Beebe, 2005; Atkinson, 1985; Halliday, 1989) and summarize four prominent sub-styles: orality, interactivity, vividness, and emotionality, supported by manual annotation. Since we aim to quantify the style elements at various positions within the stylized text and LLMs have demonstrated strong capabilities in the sentence level TST (Lai et al., 2023), for each sub-style, we generate a series of examples varying in style strength from a single sentence, and then we rank and score them with GPT3.5-Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo. Then we train a small scorer (TinyLlama-1.1b (Zhang et al., 2024)) to predict these scores.
@@
For semantic preservation evaluation, we propose a QA-based method. The accuracy changes of a QA model (eg. Llama3-8B-Instruct (AI@Meta, 2024)) on texts before and after style transfer reflect the LLM’s ability of semantic preservation. Specifically, we use GPT-4333https://openai.com/research/gpt-4 to generate high-quality QA pairs from the source text focusing on two dimensions: key information and logical structure (Coopman and Lull, 2018).
@@
  • We collect data and introduce a valuable and extensible task named Public-Speaking Style Transfer (Section 3). We decompose the intricate text style into key sub-styles for more accurate style definition and evaluation (Section 3.3 and 3.4).
@@
Table 1: Comparison of PSST and formality TST: a. “p-s” means public-speaking style. b. Traditional TST primarily focuses on vocabulary-level adjustments, such as adhering to writing norms and word norms. In contrast, PSST involves audience engagement, including posing questions to generate interest and appeal to the audience.
@@ 2 Related Work
Style, from a linguistic perspective, encompasses various elements that contribute to the conveyance of semantics, including word choice, sentence structure, and arrangement, all of which work together to establish the tone, imagery, and meaning in the text (McDonald and Pustejovsky, 1985; Hu et al., 2022; Li et al., 2024b). In contrast, research on text style transfer (TST) takes a data-driven approach, defining style as attributes or labels based on style-specific corpora  (Shen et al., 2017b; Rao and Tetreault, 2018), which may affected by other attributes in dataset (Jin et al., 2022). However, our approach is grounded in real data and explores key dimensions from the perspective of public-speaking linguistics. This definition alleviates the ambiguity and reduces the difficulty of data construction and evaluation.
@@
Firstly, fluency, a common objective in most natural language generation tasks, is often measured by the perplexity score (PPL) (Yang et al., 2018). Secondly, to evaluate content preservation during the style transfer, metrics include BLEU (Papineni et al., 2002), ROUGE (Lin and Och, 2004), BERTScore (Zhang et al., 2019) are employed. Thirdly, the strength of style is an important dimension. Typically, a binary style classifier is first separately pretrained to predict the style label of input sentences (Gao et al., 2024; Lai et al., 2021; Zhan et al., 2022). This classifier is then used to estimate the style transfer accuracy. Recently,  Lai et al. (2023) demonstrates that ChatGPT (OpenAI, 2022) achieves competitive correlation with human judgments to serve as a multidimensional evaluator for sentence-level formal-informal style transfer.
@@ 3.4 Key Features of Public-Speaking Style
  • Interactivity: Interactivity in public speaking refers to the speaker engaging with the audience through various means such as posing thought-provoking, facilitating personal reflection, and crafting intriguing hypothetical scenarios (Table 1 example 1).
  • Emotionality: Public speaking contains the speaker’s appropriate views and attitudes on specific events to reflect the speaker’s emotional tendencies and inner thoughts (Table 1 example 2).
  • Vividness: In public speaking, speakers should present information in a lively, easy-to-understand way, such as using analogies and metaphors to make complex ideas more accessible and engaging (Table 1 example 3).
  • Orality: The public-speaking style texts should align with oral communication norms, incorporating appropriate filler words, simple sentence structures, and suitable word choices (Table 1 example 4)
@@ 4.1 Style Strength Evaluation
To evaluate the style strength of long texts in the PSST task, we propose a multi-dimensional, fine-grained evaluation method. Specifically, we generate a series of examples with varying style strengths from a single sentence, rank and score them using gpt-3.5-turbo, and then train a small scorer (TinyLlama-1.1b (Zhang et al., 2024)) to predict these scores. This approach is appropriate for three main reasons: (1) Early experiments suggest that using LLMs directly as evaluators may not be feasible (Section 4.1.3). (2) Relevant studies show that LLMs possess strong stylization and evaluation capabilities in TST tasks at the sentence level (Lai et al., 2023). (3) For long texts, a simple overall style strength score does not facilitate detailed analysis. We believe that a similar style of two long texts should be reflected in the overall style strength and the stylization of specific positions within the documents, such as the beginning and the end.
@@ 4.2.2 Correlation
We conduct three manual evaluation experiments to verify the effectiveness of the QA-based method.
@@
The experiments above demonstrate that the decreased accuracy of the QA model is indeed attributed to the semantic loss occurring during the style transfer process. Please refer to Appendix E.3 for more details.
@@ 5.1 Experiment Settings
We assessed the public-speaking style modeling capability of the LLM by comparing it with the following text. (1) src_text, the official texts mentioned in Section 3.2, characterized by an official style and complete information, serving as the upper bound for semantic preservation. (2) paraphrase, texts paraphrased by gpt-3.5-turbo based on the source text, reflecting inherent changes in the language processing of LLMs. (3) ted_text, the target style text in real public speaking scenario. Note that src_text and ted_text are not parallel. Both src_text and paraphrase have low public-speaking style strength but high semantic preservation scores. The closer the style strength of the generated texts is to real-world data and the lower their semantic loss, the stronger the model’s capability in language style modeling.
@@
(1) concise prompt, a brief instruction requiring models to simulate lively and engaging public speaking through style transfer. (2) enhanced prompt, detailed guidance on the dimensions of stylization, cautioning against excessive stylization, and emphasizing the importance of semantic consistency. Five prompts are created for each type shown in Figure 16.
@@
We employ Llama 2-Chat (Touvron et al., 2023), Llama 3-Instruct (AI@Meta, 2024) and gpt-3.5-turbo to generate public-speaking style texts based on src_text and prompts. We believe these models capture the current state of LLMs and can help explore the gap between LLMs and human abilities in complex, abstract language style modeling.
@@ 5.2 Evaluation Results and Analysis
The results comparing style strength distribution across different positions within texts are illustrated in Figure 5. Specifically, in terms of interactivity and orality, the style distribution in real-world scenarios is more uniform, with slight increases in stylization at the opening and closing segments. This pattern aligns with the conventional use of introductory and concluding phrases in real-world settings. However, for LLMs, the distribution exhibits a “U-shaped” pattern, indicating excessive stylization at the text’s beginning and end. For instance, as depicted in Figure 5, compared to real-world texts, LLMs such as GPT-3.5, Llama-3-70B-Instruct, and Llama-2-Chat-70B demonstrate pronounced stylization at the beginning and end, but less so in the middle sections (Case in Figure 21). This disparity suggests that LLMs diverge from human-like style modeling, potentially due to a model-inherent tendency to focus on the beginning and end of long texts, or possibly indicative of the model’s “lazy” behavior.
@@
The results shown in Figure 6 reveal significant semantic loss during style transfer. The high accuracy observed in src_text and paraphrase confirms that the QA model can effectively respond to questions based on non-stylized texts. Moreover, our questions, designed to target essential information, are relatively straightforward and can be answered provided no information is lost. Therefore, the observed decline in accuracy with stylized texts further underscores the challenges LLMs face in semantic preservation. As shown in Figure 24, the stylized text omits detailed descriptions of key steps, replacing them with generalized and vague statements, which is unfavorable, particularly in public education contexts such as popular science communication. In Figure 25, the stylized text lacks a temporal reference for the specific event, which impairs the comprehensive representation of the event, particularly in the dissemination of critical information, such as news reporting. Notably, models exhibiting stronger stylization (e.g., LLaMA-3-70B-Instruct vs. GPT-3.5-Turbo) tend to obscure key information more. This is intuitive, as excessive stylization reduces the clarity of semantic content, thus hindering message comprehension for audiences, even if the essential points remain present. Additionally, while enhanced prompts focus on semantic preservation and achieve higher accuracy, they are still less than satisfactory.
@@ 6 Conclusion
In this paper, we introduce public-speaking style transfer, which requires LLMs to transform a formal and long text into a public-speaking style. By analyzing real-world data from a linguistic perspective, decompose the intricate text style into key sub-styles. We propose a fine-grained evaluation framework that enables an evaluation-driven approach to continuously assess and enhance the language style modeling capabilities of LLMs. For style strength evaluation, we distill the LLM’s ability to assess the style elements of sentences into a smaller model for each sub-style. We introduce two metrics: passage-level style strength score and style strength distribution, allowing for a detailed analysis of long stylized texts. For semantic preservation, we propose a QA-based method that focuses on key information and logical structure. By observing changes in the QA model’s accuracy, we identify and analyze the LLM’s ability to retain semantics. Our experiments reveal that current LLMs exhibit issues such as over-stylization, uneven style strength distribution, and severe semantic degradation in long and complicated language style modeling. These findings highlight the shortcomings in LLMs’ language style modeling capabilities and underscore the substantial room for improvement.
@@
Token Length and Variety of Tested Models
Limited Token Length
Given the substantial semantic losses observed in current LLMs, we advocate for studying their complex style modeling capabilities within bounds that are reasonable given their current performance levels.
Limited Variety of Models Tested
We primarily focused on testing mainstream, widely used, and capable LLMs (such as LLaMA and GPT-3.5). We believe these models are representative of the current landscape of LLMs and can effectively explore the gap between current LLMs and human abilities in complex, abstract language style modeling.
In future work, we plan to investigate longer texts and a broader range of models to further validate and demonstrate the generalizability of our findings.
@@ Acknowledgements
We thank Xinyue Liang and Shunyu Wang for their support in human annotation for this study.
@@ A.1 Target Dataset: Public Speaking Style Data from Real Scenarios
It is important to note that we utilize the above data to analyze general stylistic characteristics. Furthermore, in this paper, we primarily employ TED data as a representative example to demonstrate the effectiveness of our evaluation pipeline. Likewise, the same evaluation pipeline can be applied to data from other scenarios to achieve comparable results.
@@ C.1 Results of Different Text Lengths in PSST
Table 7: Style Strength Evaluation Results($\textbf{800}\pm 200$ tokens). a. The original score ranges from 0 to 5. b. src_text means Source Text and paraphrase means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).
@@
Table 8: Style Strength Evaluation Results($\textbf{400}\pm 100$ tokens). a. The original score ranges from 0 to 5. b. src_text means Source Text and paraphrase means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).
@@
Table 10: Style Strength Evaluation Results($\textbf{1200}\pm 100$ tokens). a. The original score ranges from 0 to 5. b. src_text means Source Text and paraphrase means Paraphrased Text, both are styleless. c. Scored by TinyLlama-1.1b(Reward).